# Collaborating with AI Agents: Field Experiments on Teamwork, Productivity, and Performance

### v1

@@ 3.2 Productivity
Human-AI teams are near-substitutable with Human-Human teams with half as many individuals.
On the MindMeld platform, participants were free to submit as many ads as they could produce within the time limit. On the team level, productivity was comparable between Human-Human and Human-AI teams (Table 5). However, on the individual level, participants in Human-AI teams individually submitted 60% to 73% as many ads as their counterparts in Human-Human teams. This suggests that AI collaboration supports individual productivity by enabling participants to generate more outputs compared to those in Human-Human teams.
@@
This focus on individual task completion is further reflected in ad copy completion rates. While completing tasks may seem like a minimal benchmark for productivity, it is particularly relevant for participants with lower performance levels, such as workers on platforms like Prolific. As shown in Table 6, with model-free evidence shown in Figure 7, participants in Human-AI teams had consistently higher completion rates for ad copy elements (e.g., headline, primary text, description) compared to participants in Human-Human teams. This finding suggests that AI collaboration provides support for individuals who might otherwise struggle to complete tasks, aligning with prior research on AI’s role in enhancing performance for lower-performing participants.
@@
3.3 Performance
We now evaluate the performance of Human-Human and Human-AI teams by assessing the quality of their outputs. Specifically, we examine differences in the quality of ad text and images and field evaluations of ad effectiveness to understand how collaboration dynamics influence overall effectiveness.
@@ 3.3 Performance
To determine the quality of ads, we use two post hoc measurements of ad quality: human evaluations and AI evaluations. For human evaluations, we recruited 1,195 participants from Prolific to rate the text quality, the image quality, and the estimated likelihood of clicking on the ad (see Section 2.9 for details). The participants were shown a mockup of each ad as it would appear on a social media platform (Figure 3). In our AI evaluations, we used gpt-4o-mini to ask the AI model the same questions we ask the human evaluators (see Section 2.8 for details).
In our human evaluations of ad quality, we found that Human-AI teams had higher-quality text but lower-quality images than Human-Human teams (Table 7). We show model-free evidence in Figure 8. The estimated likelihood of clicking on the ad was indistinguishable between the two groups. Interestingly, in our AI evaluations of ad quality, we found that the AI ratings were higher on the text and clicks for ads produced by Human-AI teams and the same across the groups for image quality. In a way, it is unsurprising that the AI rated these ads’ text quality, image quality, and estimated click likelihood as equal to or better than those produced by Human-Human teams because the ads by Human-AI teams were created with the assistance of OpenAI’s gpt-4o. However, human ratings show that AI introduces trade-offs: collaborating with GPT models improves text quality but reduces image quality. This trade-off is perhaps because GPT models were designed primarily for next-word prediction vs predicting image quality. These results suggest that while GPT models enhance text-focused tasks, their contributions to multimodal outputs like ads may require complementary tools or models designed specifically for image-related tasks. These findings also raise the question of whether and how these trade-offs in text and image quality meaningfully impact actual clickthrough rates.
@@
Sex[Male] 0.078∗∗∗ 0.000 0.026 0.016 0.003 -0.008
(0.017) (0.017) (0.018) (0.014) (0.014) (0.012)
Demographics Yes Yes Yes Yes Yes Yes
Observations 11138 11138 11138 11138 11138 11138
Table 7: AI and Human ratings of ads.
@@
3.4 Personality prompts
The second pre-registered set of randomization in our study focuses on AI personality traits. Specifically, we manipulate the Big Five personality traits for each AI, independently setting them to high or low levels using P2 prompting (Jiang et al., 2023). This allows us to systematically investigate how AI personality traits influence collaborative work and whether there is heterogeneity in their effects based on the personality traits of the human collaborators, as measured through a pre-task survey. For more details on this approach, see Section 2.1.
Teamwork.
Examining the interaction effects of prompts with human personalities reveals significant heterogeneous prompt effects on collaboration outcomes (Table LABEL:tb:counts:personality:teamwork). First, on the level of communication measured by the number of messages, we see that a conscientious AI increases communication by a substantial amount, 62%, but only if the human counterpart is also high on conscientiousness. This intuitive result indicates a synergy between the conscientiousness of the AI and that of the human counterpart, where the alignment of traits, and not the traits individually, increases communication.
Table 8: Personality-heterogeneous prompt effects on collaboration outcomes on the individual level.
Messages Copy Edits Image Selects AI-Generated Submissions
Images
Intercept 28.033 -417.018 41.960 5.430 5.356
(15.798) (333.169) (16.944) (3.869) (3.014)
OpennessAI 4.629 68.418 -1.154 -0.538 -0.329
(12.011) (264.775) (13.691) (3.166) (2.096)
[…]
(7.145) (165.372) (7.789) (1.575) (1.224)
[…]
(8.950) (197.030) (9.685) (2.213) (2.090)
[…]
(7.294) (149.223) (8.502) (1.789) (1.315)
[…]
(7.951) (179.521) (8.084) (1.701) (1.676)
[…]
(6.918) (164.836) (9.042) (1.940) (1.264)
ConscientiousnessAI -8.670 277.472 10.499 -0.912 0.149
(11.251) (260.988) (14.708) (3.102) (2.433)
[…]
(6.861) (165.094) (7.629) (1.571) (1.265)
[…]
(8.408) (194.647) (9.875) (2.177) (2.302)
[…]
(7.318) (159.277) (8.211) (1.764) (1.314)
[…]
(7.475) (198.311) (8.107) (1.676) (1.442)
[…]
(6.636) (159.721) (9.050) (1.854) (1.343)
ExtraversionAI -12.432 431.725 -18.305 -2.513 0.390
(11.693) (276.559) (14.277) (3.242) (2.303)
[…]
(7.059) (157.808) (7.788) (1.627) (1.265)
[…]
(8.731) (208.164) (10.336) (2.319) (2.207)
[…]
(7.330) (155.943) (7.779) (1.725) (1.298)
[…]
(8.016) (181.480) (8.002) (1.645) (1.444)
[…]
(7.078) (164.950) (8.726) (1.875) (1.237)
AgreeablenessAI 0.741 270.288 0.796 2.037 -0.428
(11.380) (269.627) (13.630) (3.210) (2.259)
[…]
(7.018) (162.741) (7.579) (1.618) (1.286)
[…]
(7.969) (188.352) (9.415) (2.196) (1.976)
[…]
(7.430) (152.002) (7.852) (1.758) (1.305)
[…]
(7.863) (182.273) (8.048) (1.569) (1.360)
[…]
(6.710) (163.766) (9.180) (1.886) (1.228)
NeuroticismAI 4.905 690.342∗∗ -14.471 3.209 3.066
(12.325) (260.106) (14.277) (3.329) (2.267)
[…]
(7.055) (159.805) (7.701) (1.651) (1.302)
[…]
(8.786) (189.858) (9.740) (2.229) (2.235)
[…]
(7.252) (154.798) (8.068) (1.736) (1.305)
[…]
(8.303) (183.514) (8.440) (1.803) (1.683)
[…]
(7.148) (161.053) (9.015) (1.966) (1.320)
OpennessH 14.867 393.999 -0.717 1.067 2.404
(8.956) (201.678) (9.834) (1.945) (1.519)
ConscientiousnessH -8.065 77.226 1.701 -0.158 1.052
(11.145) (263.025) (12.535) (2.768) (2.722)
ExtraversionH 11.162 -41.462 -9.888 0.928 -1.894
(9.525) (189.557) (9.505) (2.163) (1.680)
AgreeablenessH -11.982 554.955 -13.027 0.031 0.085
(10.785) (231.536) (11.147) (2.230) (2.183)
NeuroticismH 3.706 345.464 6.556 0.412 -2.405
(9.506) (229.219) (10.524) (2.197) (1.810)
Demographics (+ interaction terms) Yes Yes Yes Yes Yes
Observations 1258 1258 1258 1258 1258
Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Heteroskedasticity-robust standard errors are reported. AI{AI} refers to induced AI personality; $H$ refers to human personality covariates.
Table 8: *
Next, we observed that AI neuroticism significantly influences edits on the display ad, including both copy and image selection. While a neurotic AI increased the number of copy edits, this effect reversed when the AI was paired with an agreeable human. For image selection, where participants choose from a set of images, a neurotic AI only increased the number of selections when collaborating with an open human. These findings underscore the nuanced ways in which AI and human personality traits interact to shape distinct facets of collaborative ad creation, such as editing and image selection.
Productivity.
For productivity, we observed that AI agreeableness had significant heterogeneous effects on the number of submissions, on the individual level. While an agreeable AI by itself does not have any main effects, it increases the number of submissions by 2-3 ads when paired with either an extraverted or neurotic human. This highlights that distinct patterns also emerge for productivity outcomes.
Performance.
To understand the effect of personality prompts on performance, we estimated heterogeneous prompt effects on the quality of completed ads. First, we examined the effects on our post hoc human evaluations of ad quality. In our analyses shown in Table LABEL:tb:counts:personality:performance, we observed numerous significant prompt effects, many of which have opposite heterogeneous effects depending on the personality of the participant. Interestingly, for text quality, we observed that a conscientious AI has a negative overall effect on text quality. However, this negative effect is reversed when the human is conscientious or open. AI extraversion has a negative effect when the human is agreeable but a positive effect when the human is extraverted. AI agreeableness has an overall positive effect, but this is reversed when the human is open or conscientious. AI neuroticism has a negative effect when the human is extraverted or neurotic but a positive effect when the human is agreeable.
Table 9: Personality-heterogeneous prompt effects on quality outcomes on the individual level.
Human evaluations
Text Image Click
Intercept 4.856∗∗∗ 4.434∗∗∗ 3.342∗∗∗
(0.232) (0.245) (0.260)
OpennessAI 0.155 -0.014 0.214
(0.174) (0.184) (0.193)
OpennessAI $\times$ OpennessH 0.184 -0.216 -0.052
(0.105) (0.111) (0.117)
OpennessAI $\times$ ConscientiousnessH -0.073 0.407∗∗ 0.060
(0.130) (0.142) (0.149)
OpennessAI $\times$ ExtraversionH 0.112 0.183 0.259
(0.111) (0.118) (0.123)
OpennessAI $\times$ AgreeablenessH 0.039 -0.341∗∗ -0.144
(0.105) (0.111) (0.115)
OpennessAI $\times$ NeuroticismH -0.069 0.008 -0.143
(0.102) (0.106) (0.111)
ConscientiousnessAI -0.628∗∗∗ 0.134 0.046
(0.171) (0.182) (0.190)
ConscientiousnessAI $\times$ ConscientiousnessH 0.392∗∗ 0.075 0.082
(0.128) (0.139) (0.144)
ConscientiousnessAI $\times$ ExtraversionH -0.406∗∗∗ -0.449∗∗∗ -0.407∗∗
(0.110) (0.118) (0.123)
ConscientiousnessAI $\times$ OpennessH 0.261 0.104 0.145
(0.104) (0.111) (0.116)
ConscientiousnessAI $\times$ NeuroticismH -0.168 -0.167 -0.278
(0.099) (0.104) (0.108)
ConscientiousnessAI $\times$ AgreeablenessH 0.047 -0.100 -0.097
(0.105) (0.109) (0.113)
ExtraversionAI 0.232 0.139 0.103
(0.173) (0.185) (0.193)
ExtraversionAI $\times$ OpennessH -0.072 -0.142 -0.157
(0.103) (0.111) (0.116)
ExtraversionAI $\times$ AgreeablenessH -0.423∗∗∗ -0.073 -0.210
(0.106) (0.111) (0.115)
ExtraversionAI $\times$ ExtraversionH 0.376∗∗ -0.109 0.161
(0.109) (0.117) (0.123)
ExtraversionAI $\times$ ConscientiousnessH -0.172 0.040 0.030
(0.133) (0.143) (0.148)
ExtraversionAI $\times$ NeuroticismH 0.134 0.080 0.169
(0.102) (0.107) (0.112)
AgreeablenessAI 0.491∗∗ -0.158 -0.101
(0.168) (0.176) (0.184)
AgreeablenessAI $\times$ OpennessH -0.437∗∗∗ 0.237 0.123
(0.102) (0.108) (0.113)
AgreeablenessAI $\times$ ConscientiousnessH -0.597∗∗∗ -0.327 -0.369
(0.133) (0.141) (0.148)
AgreeablenessAI $\times$ ExtraversionH 0.173 -0.001 0.108
(0.110) (0.117) (0.123)
AgreeablenessAI $\times$ AgreeablenessH -0.071 0.022 0.093
(0.106) (0.111) (0.116)
AgreeablenessAI $\times$ NeuroticismH 0.122 -0.070 0.030
(0.101) (0.105) (0.109)
NeuroticismAI 0.068 0.210 0.168
(0.175) (0.185) (0.195)
NeuroticismAI $\times$ OpennessH 0.008 0.046 0.022
(0.103) (0.109) (0.115)
NeuroticismAI $\times$ ConscientiousnessH 0.105 -0.004 0.111
(0.130) (0.139) (0.145)
NeuroticismAI $\times$ ExtraversionH -0.267 0.094 -0.008
(0.110) (0.117) (0.123)
NeuroticismAI $\times$ AgreeablenessH 0.356∗∗ -0.055 -0.050
(0.107) (0.112) (0.117)
NeuroticismAI $\times$ NeuroticismH -0.236 -0.223 -0.095
(0.104) (0.107) (0.112)
OpennessH 0.161 0.088 0.039
(0.127) (0.136) (0.145)
ConscientiousnessH 0.154 -0.055 0.111
(0.173) (0.188) (0.197)
ExtraversionH -0.210 0.063 -0.122
(0.130) (0.141) (0.149)
AgreeablenessH 0.078 0.289 0.272
(0.142) (0.151) (0.159)
NeuroticismH 0.181 0.240 0.263
(0.136) (0.144) (0.152)
Demographics (+ interaction terms) Yes Yes Yes
Observations 29874 29874 29874
Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Heteroskedasticity-robust standard errors are reported. Units are on a 7-point Likert scale. AI{AI} refers to induced AI personality; $H$ refers to human personality covariates.
Table 9: *
Next, we examined the heterogeneous effects of AI prompts on image quality. These effects are particularly important, given that ads created by Human-AI teams were rated as having lower image quality compared to those from Human-Human teams. For conscientious humans, AI openness improved image quality, while for agreeable humans, it had the opposite effect. AI conscientiousness negatively impacted image quality, but only when paired with extraverted humans. In contrast, AI agreeableness enhanced image quality for open humans but diminished it for conscientious humans. Finally, AI neuroticism reduced image quality specifically for neurotic humans.
Then, we analyzed the likelihood of clicking on the ads as an indicator of ad engagement. We only found that AI openness paired with human extraversion increased the click quality. Other combinations of personalities decreased click quality: AI conscientiousness paired with human extraversion or agreeableness and AI agreeableness paired with human conscientiousness, which is, interestingly, symmetrical between human and AI pairs.
3.5 Field study
To evaluate the real-world performance of ads created by Human-Human and Human-AI teams, we conducted a field experiment on a social media platform. Our ad campaigns generated 4,932,373 impressions and 7,546 clicks over 20 days. This section examines how collaboration type influences key advertising metrics: cost-per-click (CPC) and click-through rate (CTR), reported in Table 10, followed by view-through rate (VTR) and view-through duration (VTD) in Table 11. The field study extends lab findings and tests how the distinct productivity and quality profiles of Human-AI and Human-Human teams translate into advertising outcomes in a live setting. Broadly, we found that ads created by Human-AI teams performed similarly to those by Human-Human teams.
@@ 3.5 Field study
We examined CPC and CTR using regression models with campaign random effects to account for unobserved heterogeneity across the 400 campaigns, with results in Table 10. For CPC, measured in dollars, our analysis in Column 1 reveals no significant effect of collaboration type. [^6] Higher spend consistently lowers costs and suggest divergent targeting and optimization, likely because we had kept on auto-bidding in our campaigns. Thus, we consider the effects present in this section as conditional on ad algorithms, as per recommendation 5 from Braun et al. (2024). With human-evaluated quality scores (i.e., Image, Text, Click) as covariates, ads with stronger image quality significantly reduce CPC, and this pattern holds when all predictors are combined. For CTR, expressed as a percentage, the analysis with collaboration indicates no direct Human-AI effect. Adding quality scores highlights text quality as a key driver of higher click-through rates, which remains consistent when all factors are included. These results suggest that Human-AI ads perform broadly equivalent to Human-Human ads in click metrics, with outcomes shaped by quality rather than team type.
@@
We assessed view metrics—view-through rate (VTR) and view-through duration (VTR; in log-seconds)—using similar regression models, as shown in Table 11. Our analysis of VTR shows no significant effects for collaboration type or quality. Our analysis of VTD reveals no Human-AI effect but uncovers text quality as a significant driver of longer viewing duration. These findings affirm the broad equivalence of Human-AI and Human-Human ads in view-through performance. While VTR remains largely unaffected by team type or quality, our analysis of VTD shows that the higher text quality of Human-AI teams can increase engagement time as well as CTR. Overall, these results confirm that Human-AI collaboration yields ads comparable to Human-Human efforts, with text and image quality differentially shaping real-world performance.
@@
Observations 1859 1859 1859 2000 2000 2000
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001.
Personality prompts
As before in Section 3.4, we examined the interaction effects of personality prompts on the outcomes of the field study. Table 12 shows the regression terms for the significant interaction effects, and we report the full regression in Section D. Among the significant interaction terms, we found that conscientious humans had 0.088% higher CTR but 0.072% lower CTR if the AI was prompted to be neurotic. An agreeable-prompted AI reduced CPC by \$4.78, but this effect was reversed for an extraverted human by \$3.11. Finally, a neurotic human paired with a neurotic AI saw CTR improved by 0.059%. These findings underscore that personality interactions between humans and AI agents significantly shape field performance and highlight the potential for tailored AI agents to improve advertising outcomes.
CPC (\$) CTR (%)
Intercept 14.211*** 16.533*** -0.067 -0.141*
(2.920) (3.041) (0.057) (0.059)
Spend -0.144*** -0.150*** 0.007*** 0.008***
(0.019) (0.020) (0.000) (0.000)
Image -0.461** 0.005
(0.178) (0.003)
Text -0.314 0.012***
(0.169) (0.003)
Click 0.470* -0.006
(0.198) (0.004)
ConscientiousnessH -1.797 -1.913 0.088* 0.084
(2.300) (2.295) (0.045) (0.045)
AgreeablenessAI -4.779* -4.457 0.027 0.013
(2.311) (2.312) (0.045) (0.045)
ConscientiousnessH $\times$ NeuroticismAI 3.531* 3.537* -0.072* -0.068*
(1.764) (1.758) (0.034) (0.034)
ExtraversionH $\times$ AgreeablenessAI 2.991 3.105* -0.019 -0.020
(1.542) (1.537) (0.030) (0.029)
NeuroticismH $\times$ NeuroticismAI 0.956 1.156 0.059* 0.056*
(1.443) (1.439) (0.028) (0.028)
Campaign RE Yes Yes Yes Yes
Observations 1859 1859 2000 2000
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Campaign RE represents campaign random effects. Ads with zero clicks were removed for regressions for columns 1-2.
Table 12: Significant interaction effects of human and prompted AI personalities on cost-per-click (CPC) and click-through rates (CTR) from the field study.

### v2

@@ 3.2 Productivity
Human-AI teams produce a similar amount of output with half as many workers.
Participants were incentivized to submit as many ads as possible within the time limit. At the team level, productivity was comparable between human-human and human-AI teams as human-AI teams submitted approximately the same number of ads with half as many workers (Table 5). At the individual level, participants in human-AI teams submitted 73% more ads than their counterparts in human-human teams, with results consistent when demographic controls are added. AI collaboration thus supports individual productivity by enabling participants to generate more output per worker compared to those in human-human teams.
Although our findings suggest that human-AI teams produce a similar amount of output as human-human teams with half as many workers, it is important to note that our study did not include a human-alone condition. Prior studies have consistently shown that AI tools enhance individual productivity compared to humans working alone across various tasks (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024, Wiles et al., 2023). In light of this evidence, our results suggest that the AI in human-AI teams acts as a near-substitute for an additional human collaborator, enabling a single human to achieve productivity levels comparable to a team with two human workers. Without a human-alone condition, however, we cannot directly quantify AI’s incremental contribution beyond what an individual human might have achieved working alone. We explore this limitation further in the Discussion section.
@@
With that caveat in mind, to test the robustness of our results to the alternative hypothesis that individual-level productivity differences are simply due to the submission of incomplete ads, we examined ad copy completion rates across conditions. As shown in Table 6 and Figure 7, participants in human-AI teams had consistently higher completion rates for ad copy elements compared to participants in human-human teams. Taken together, these findings suggest that human-AI teams are indeed more productive, in both the number of ads submitted and the ad completion rate.
@@
3.3 Ad Quality
In addition to the productivity and work process implications of collaborating with AI agents, we also evaluated the performance of human-human and human-AI teams by assessing the quality of their output. Specifically, we examined differences in the quality of ad text and images and conducted field evaluations of ad effectiveness to understand how human-AI collaboration dynamics influence overall effectiveness.
@@ 3.3 Ad Quality
In human evaluations of the quality of the ads, we found that human-AI teams had higher quality ad text but lower quality ad images compared to human-human teams (Table 7, Figure 8). The estimated likelihood of clicking on the ad was indistinguishable between the two groups. Interestingly, in our AI evaluations of ad quality, we found that AI ratings were higher on the text and clicks for ads produced by human-AI teams and the same across the groups for image quality. In a way, it is unsurprising that the AI rated these ads’ text quality, image quality, and estimated click likelihood as equal to or better than those produced by human-human teams because these ads were created in collaboration with OpenAI’s gpt-4o.
However, the human ratings reveal that AI introduces important trade-offs in output quality. Specifically, collaborating with GPT models improves text quality but reduces image quality. This trade-off may arise because GPT models were originally designed for next-word prediction rather than to predict image quality. These results suggest that while GPT models enhance text-based outputs, their contributions to multimodal outputs like ad images may require complementary tools or models designed specifically for image-related tasks. These findings then beg the question of whether and how these trade-offs in text and image quality meaningfully impact actual click through rates, which we examined in detail in the field test of the ads themselves.
@@
Demographics Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Observations 45803 45803 45803 45803 45803 45803 11138 11138 11138 11138 11138 11138
  • Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Standard errors are clustered at the ad level to account for within-ad correlation. “H-AI” refers to human-AI teams. “Human AI Images Generated” refers to the AI images generated by humans. Copy and image edits are per thousand.
Table 7: Human and AI ratings of ads.
@@
We also examined how work process variables, like the number of messages exchanged, the number of ads submitted, the number of AI images generated, and the number of copy and image edits, affect ad quality in post hoc analyses. We found that more messages exchanged, fewer ads submitted, and more AI images generated were associated with higher human- and AI-evaluated ad quality. Interestingly, for text copy and image edits, we observed nuanced differences in ad quality across team types and human and AI evaluations. Our analyses showed that humans rate ads more highly when humans copy edit text more in human-human teams. In contrast, AI rate ads more highly when humans copy edit text less in human-human teams. Human text copy edits made in human-AI teams mostly revealed no difference in preference in either human or AI evaluations of ad quality (except for a slight negative preference for AI-evaluated click quality). Both humans and the AI rated ads higher when image edits were made in human-AI teams and rated ads lower when image edits were made in human-human teams. This finding is consistent with previous our findings that humans are better at selecting images—as human edits to images are more highly valued, by both humans and AI, in teams in which AI is contributing to the image selection and generation process, humans seem to be needed to correct, adjust or improve image selection and generation when AI is involved. These findings highlight that team composition (human-human vs human-AI teams) and work processes (messages exchanged, ads submitted, AI images generated, and copy and image edits) all meaningfully shape ad quality.
3.4 Ad Effectiveness
To evaluate the real-world performance of ads created by human-human and human-AI teams, we conducted a field experiment on the social media platform X. Our ad campaigns generated 4,932,373 impressions and 7,546 clicks over 20 days. This section examines how collaboration type influences key advertising metrics including the cost-per-click (CPC) and click-through rate (CTR), reported in Table 8, and the view-through rate (VTR) and view-through duration (VTD) shown in Table 9. The field study extends lab findings and tests how the distinct productivity and quality profiles of human-AI and human-human teams translate to advertising outcomes in a real world setting. Broadly, we found that ads with higher text quality, created more often by human-AI teams, and higher image quality, created more often by human-human teams, performed significantly better on cost per click and click through rate metrics. As human-AI teams produced higher quality text and lower quality images, their effects counteracted one another and human-AI ads performed similarly to those created by human-human teams overall.
@@ 3.4 Ad Effectiveness
We examined CPC and CTR using regression models with campaign random effects to account for unobserved heterogeneity across the 400 campaigns, with results in Table 8. For CPC, measured in dollars, our analysis in Column 1 reveals no significant effect of collaboration type. [^6] Higher spend consistently lowers costs and suggest divergent targeting and optimization, likely because we had kept on auto-bidding in our campaigns. Thus, we consider the effects present in this section as conditional on ad algorithms, as per recommendation 5 from Braun et al. (2024). With human-evaluated quality scores (i.e., Image, Text, Click) as covariates, ads with stronger image quality significantly reduced CPC by \$0.268 or 2.3% (Column 2), and this pattern held when all predictors were included in the analysis (Column 3). For CTR, expressed as a percentage, the analysis with collaboration indicated no direct human-AI effect (Column 4). With quality scores, we found that text quality significantly reduced CTR by 0.008% (Column 5), which remained consistent when all factors were included (Column 6).
@@
We assessed view metrics—view-through rate (VTR) and view-through duration (VTD; in log-seconds)—using similar regression models, as shown in Table 9. Our analysis of VTD followed a similar pattern as observed in the click metrics—while there was no direct human-AI effect (Column 4), text quality was significantly associated with longer viewing duration. On average, one point higher text quality ratings were associated with one second longer view through duration (Columns 5-6). Our analysis of VTR showed no significant effects for collaboration type or quality (Columns 1-3). These findings demonstrate that ads with higher text quality, which are more often produced by human-AI teams, are viewed for longer periods of time.
@@
Observations 2000 2000 2000 2000 2000 2000
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Campaign RE represents campaign random effects.
Table 9: Effects on view-through rate (VTR) and view-through duration (VTD) from the field study.